OpenStudy (anonymous):

If ƒ (x ) = 3x2 - x, find ƒ (-2).

hero (hero):

For example f(-3) = 3(-3)^2 - (-3) = 3(9) + 3 = 27 + 3 = 30

OpenStudy (anonymous):

If ƒ (x ) = 1/3x2 - x, find ƒ (-2).

OpenStudy (anonymous):

So in this case they're asking to plug in -2 for x because they're implying that x = -2.

hero (hero):

@junae43, is it \[f(x) = \dfrac{1}{3x^2 - x}\] or \[f(x) = \dfrac{1}{3x^2} - x\]

OpenStudy (anonymous):

thanks

hero (hero):

I'm asking you which f(x) are you working on?